teacher jane randomly selected the grades in statistics and probability of 7 grade 11 students. calculate the mean grade, variance and standard deviation of the given set of data.



1
Expert's answer
2021-05-31T15:31:29-0400
"\def\arraystretch{1.5}
 \begin{array}{c:c:c}
 x & x-\bar{x} & (x-\bar{x})^2 \\ \hline
94 & 8 & 64 \\
88 & 2 & 4\\
79 & -7 & 49 \\
89 & 3 & 9 \\
91 & 5 & 25 \\
84 & -2 & 4 \\
77 & -9 & 81 \\
 \hdashline
 \sum_i x_i=602 & & \sum_i (x_i-\bar{x})^2=236 \\ 
\end{array}"

"mean=\bar{x}=\dfrac{1}{n}\sum_ix_i=\dfrac{1}{7}(602)=86"

"Var(X)=s^2=\dfrac{1}{n-1}\sum_i(x_i-\bar{x})^2"

"=\dfrac{1}{7-1}(236)=\dfrac{118}{3}\approx39.333333"

"s=\sqrt{s^2}=\sqrt{\dfrac{118}{3}}\approx6.271629"